# ManipGPT: Is Affordance Segmentation by Large Vision Models Enough for Articulated Object Manipulation?

### v2

@@ V-E Quantitative Evaluation in Simulator
Following the settings from ManipLLM [5], the baseline evaluation focused on the initial movement phase, which is crucial for assessing a model’s ability to understand and respond to an object’s initial state—key for successful long distance manipulation. All models were trained on a standardized dataset of 5 objects per category across 18 categories and evaluated under the same conditions. For ManipLLM, we used its pretrained model as it was trained on a larger set of objects and affordance data than ours. As a result, the train and test categories in Table II are not directly comparable, and the results reflect ManipLLM’s generalization capability rather than performance on our dataset.
The results in Table II reveal that our approach consistently achieves high success rates, averaging 52.7% on seen and 57.3% on unseen object categories. This highlights its strong generalization capability, even for unfamiliar objects. Such success can be attributed to the effective use of visual cues, robust affordance segmentation, and the pre-trained vision transformer fine-tuned with minimal data.
The random-point policy, where the robot selects a random action point within the affordance mask, achieved a 52.8% success rate on average. However, this approach is less effective for real-world applications, as points on protruding features (e.g. handles) or near revolving joints (e.g. door hinges) often lead to poor attachment or excessive force during manipulation, reducing success rates for certain categories.
Multi-step manipulations demonstrated the highest success rates, achieving 65% on training categories and 68% on test categories. This underscores the value of impedance control in handling a variety of articulated objects, highlighting the robustness and adaptability of our algorithm.
Figure 5: We tested 11 objects in our real-world experiment, numbered 1–11 as listed in Table III. Objects 1–6 belong to the pushing category, while 7–11 are in the pushing category.
@@
Figure 6: This figure demonstrates the real object experiment. Above and below show pulling and pushing manipulation, respectively.
@@ V-F Quantitative Evaluation in Real-world
We evaluated our framework in a real-world scenario where a real robot arm interacted with actual objects. The experiment included five pulling object categories and six pushing object categories, with two objects per category, each tested three times, resulting in six repetitions per category. RGB images and depth maps were captured using a RealSense D415 camera, and our model was used to predict affordances, which were then projected into 3D space. We implemented the Franka Emika Panda robot for the experiment. Since the robot originally had a parallel finger gripper, we modified it into a sticky gripper by closing the fingers and attaching double-sided foam tape around them. This adaptation allowed the gripper to function similarly to a suction gripper, mimicking the simulation environment. Although the tape did not have sufficient adhesive strength to hold heavy objects for extended periods, it was durable enough to demonstrate our framework’s performance.
The results of real-world experiment are provided in Table III. We observe that our method shows no significant sim-to-real gap thanks to the synthetic image dataset. The model achieves a high success rate, excelling particularly in manipulating objects with well-defined boundaries and low actuation force requirements, such as boxes or keyboards. However, it occasionally struggles with objects featuring transparent or small targets like dispensers, or those with ambiguous boundaries like staplers. Additionally, accurate segmentation affordance for monitors can be challenging due to light reflections on the screen.
As we tested the simulation using simulated prompts and the real world using real image prompts, the real-world visualization in Figure 6 demonstrates that our model can successfully generalize to real-world robotic scenarios.
TABLE IV: All models were tested on real images only. ”*” indicates our model trained only on simulated images.
Model SegGPT Ours* Ours
@@ V-G Ablation Study
Our ablation study investigates the impact of including real images in the training set. We compared three configurations: SegGPT, ManipGPT trained only on SAPIEN images, and ManipGPT trained with both real and SAPIEN images.
Table IV shows that even a small number of real images significantly enhance SegGPT’s generalization ability for affordance segmentation. Specifically, including real images in the training set led to a 31.1% increase in the F1 score and a 33.2% increase in the mIoU value. Given the importance of real-world images for manipulation tasks, our results strongly demonstrate the effectiveness of incorporating real data into the training process for real-world applications.
@@ V-H Limitations and failure cases
Our framework occasionally fails to produce masks, particularly with small objects or novel viewpoints. This is likely due to limited training data or ambiguity between the prompt image and the scenario image. To mitigate this, increasing the training dataset or adopting a few-shot learning approach with multiple prompts could enhance feature learning and improve mask accuracy. Additionally, the model’s reliance on one-shot prompts limits its effectiveness for unseen categories especially when a corresponding prompt image is unavailable. Developing a method to select optimal prompts from a database based on similarity could address this limitation.
Another challenge arises with transparent objects, such as kitchen pot lids, where their transparency hinders both segmentation and normal map acquisition, leading to lower success rates, as observed with the ’Kitchen Pot’ category in Table II. One potential solution is to train models on video data, as the temporal consistency across frames could help disambiguate segmentation and provide more context for handling such objects.
Selecting optimal manipulation points also remains challenging. Although our approach targets the center of the masked region after filtering, it doesn’t always ensure the highest success rate, particularly with heavy objects and revolute joints. Using more detailed prompts with annotated manipulation points could improve accuracy.
@@ VI Conclusion
In this study, we present a comprehensive approach to robotic manipulation that integrates in-context vision with responsive robotic control. By fine-tuning a large vision model with a specialized dataset and leveraging normal vectors for interaction guidance, our system can effectively handle complex manipulation tasks. The approach simplifies the process by using a single RGB image and a pair of prompt images to directly generate affordance masks, removing the need for iterative sampling techniques. Additionally, our synthetic dataset of simulated and realistic images of articulated objects enhances the robot’s performance with minimal data. The results demonstrate significant advancements in affordance prediction, enabling the robot to identify contact areas and motion directions for precise manipulation. This streamlined framework is especially valuable for creating cost-effective, adaptable robotic systems in resource-constrained environments.

### v3

@@ V-E Quantitative Evaluation in Simulator
Following the settings from [5], the baseline evaluation focuses on the initial movement phase, which is crucial for assessing a model’s ability to understand and respond to an object’s initial state—key for successful long distance manipulation. All models are trained on a standardized dataset of 5 objects per category across 18 categories and evaluated under the same conditions. For ManipLLM, we use its pretrained model as it is trained on a larger set of objects and affordance data than ours. As a result, the train and test categories in Table II are not directly comparable, and the results reflect ManipLLM’s generalization capability rather than performance on our dataset.
The results in Table II reveal that our approach consistently achieves high success rates, averaging 53% on seen and 57% on unseen object categories. While this may seem counterintuitive, the higher success rate on unseen categories can be attributed to structural similarities. Many unseen objects share articulation mechanisms (e.g., doors, lids) with those in the training set, despite belonging to different categories. Variation within training categories may also make seen objects harder to manipulate than structurally similar yet unseen ones.
The aforementioned random-point policy achieve success rate of 52% on seen and 48% on unseen categories. However, this approach is less effective for real-world applications, as points on protruding features (e.g. handles) or near revolving joints (e.g. door hinges) often lead to poor attachment or excessive force during manipulation, reducing success rates for certain categories.
Multi-step manipulations demonstrate the highest success rates, achieving 65% on train and 68% on test categories. This underscores the value of adaptive control in handling a variety of articulated objects, highlighting the robustness and adaptability of our algorithm. Multi-step evaluations are not conducted for baselines as they already showed lower single-step success rates. Given that single-step failure rates are significant, applying these baselines to multi-step tasks would not provide meaningful insights.
Figure 5: We test 11 objects in our real-world experiment, numbered 1–11 as listed in Table III. Objects 1–6 belong to the pushing category, while 7–11 are in the pushing category.
@@
Figure 6: This figure demonstrates the real object experiment process. Above and below show pulling and pushing manipulation, respectively.
@@ V-F Quantitative Evaluation in Real-world
We evaluate our framework using a real robot interacting with physical objects. The experiment covers five pulling and six pushing object categories, with two object instances per category. Each object is tested three times, totaling six trials per category. RGB and depth images are captured using a RealSense D415, and our model predicts affordance masks, which are projected into 3D space to determine the end-effector’s target pose.
We use a Franka Emika Panda robot, originally equipped with a parallel gripper. To replicate suction-based interactions from simulation, we convert it into a sticky gripper by taping its closed fingers with double-sided foam. Although the grip isn’t strong enough to hold heavy objects long-term, it is sufficient to demonstrate manipulation performance.
The results of real-world experiment are provided in Table III. We observe that our method shows no significant sim-to-real gap thanks to the synthetic image dataset. The model achieves a high success rate, excelling particularly in manipulating objects with well-defined boundaries and low actuation force requirements, such as boxes or keyboards. However, it occasionally struggles with objects featuring transparent or small targets like dispensers, or those with ambiguous boundaries like staplers. Additionally, accurate affordance segmentation for monitors can be challenging due to light reflections on the screen.
Using real image prompts for real-world testing, as opposed to simulated prompts in simulation, our results in Figure 6 show that the model generalizes well to real-world robotic scenarios.
TABLE IV: Ablation Study. All models are tested on real images only.
Model SegGPT Ours(sim) Ours
@@ V-G Ablation Study
Our ablation study investigates the impact of including real images in the training set. We compare three configurations: SegGPT, ours trained only on simulated images, and ours trained with both real and simulated images.
Table IV shows that even a small number of real images significantly enhance SegGPT’s generalization ability for affordance segmentation. Specifically, including real images in the training set lead to a 31.1% increase in the F1 score and a 33.2% increase in the mIoU value. Given the importance of real-world images for manipulation tasks, our results strongly support the effectiveness of incorporating real data into the training process for real-world applications.
@@ V-H Limitations and failure cases
Our framework occasionally fails to produce affordance masks, particularly for small objects or novel viewpoints, likely due to limited training data or ambiguity between the prompt and query images. Expanding the dataset or adopting a few-shot learning approach with multiple prompts could mitigate this. Additionally, the reliance on one-shot prompts limits effectiveness for unseen categories when a suitable prompt is unavailable; integrating a similarity-based prompt retrieval system could address this issue.
Challenges also arise with transparent objects, such as kitchen pot lids, where transparency hinders both segmentation and normal map estimation, leading to lower success rates (e.g., Kitchen Pot category in Table II). Training on video data to leverage temporal consistency could improve segmentation robustness.
Finally, selecting optimal manipulation points remains challenging. Our approach targets the center of the mask after filtering, but this may not be ideal for heavy objects or revolute joints. Incorporating more detailed prompts with annotated manipulation points could improve precision.
@@ VI Conclusion
We present a vision-driven approach to robotic manipulation that integrates in-context affordance prediction with responsive control. By fine-tuning a large vision model and leveraging normal vectors for interaction guidance, our system efficiently identifies manipulatable regions using a single RGB image and prompt pair, eliminating the need for robot data. Experiments prove robust affordance prediction, enabling precise contact localization and motion direction estimation.
While our experiments showcase suction-based manipulation, it is adaptable to other grippers. Once the affordance predictor localizes the actionable region, one can replace the suction-oriented action proposer into a 6D grasp-pose estimator (e.g., GraspNet-1b [35]) for parallel-jaw grasping on the segmented region. The impedance control stage could remain relevant for safe, compliant manipulation.
VII Acknowledgments
We thank Dr. Xinlong Wang and Dr. Xiaosong Zhang for their technical suggestions that improved our training design. This project was supported by the National Youth Talent Support Program (8200800081), National Natural Science Foundation of China (62376006) and National Natural Science Foundation of China (62136001)